sam bought two books.one book cost 4.00$ more than three times other.together,the two books cost him 72$
Answer by macston(5194)

.
What do you want to know? Always post your question.
.
X=cost of book 1; Y=cost of book 2=3X+$4.00
.
X+Y=$72 . Substitute (3X+$4) for Y.
X+(3X+$4)=$72
4X=$68
X=$17
.
ANSWER 1: Book 1 cost $17.
.
Y=3X+$4=3($17)+$4=$51+$4=$55
.
ANSWER 2: Book 2 cost $55.
.
CHECK:
X+Y=$72
$17+$55=$72
$72=$72
.